Consider the following data. I/S 2015 2016 Sales (S) 1,200 1,320 (+10%) - Costs (C) (COGS...

  1. Consider the following data.

I/S

2015

2016

Sales (S)

1,200

1,320 (+10%)

- Costs (C) (COGS &  SG&A)

1,000

= EBITDA (=EBIT)

200

- Interest

20

=EBT

180

- Tax (T)

40

=      NI

140

  • Dividend

40

  • Plowback

100

B/S

Assets (A)

2,000

Debt (D)

800

Equity (E)

1,200

  • Common stock (CS)

           800

  • Retained earnings (RE)                

400(+100)

From these data, calculate the following ratios, showing all work:

Margin Cost =

Turnover (TO)

Interest Rate

Tax Rate

Leverage

Assume that depreciation is included in “Costs”. Using these ratios, calculate the I/S and B/S for 2016.

Assume that the leverage remains constant (at 0.40).

Assume the common stock remains constant (at 800), that is no new common stock is issued.

1. Continued

  1. Using these data, calculate the cash flow to be discounted (CF) for 2016.
  1. Assuming that this CF increases by 10% per year indefinitely, what is the value of the company? (Use a discount rate (r) of 15%)

  1. What is the dividend paid during 2016? What is the dividend yield? What is retained earnings (RE) at the end of 2016?
Answer #1

Margin Cost = $ 1000 - i.e COGS is the marginal cot of production, since we donet have the brekadown of Fixed cost or variable cost.

Turnover (TO) = Sales/ Assets = 1200/2000 = 0.60

Interest Rate = Interest is paid on Debt so interest is paid is 20 on debt 800 = 20/800 = 2.5%

Tax Rate = Tax is calculated on income before tax (EBT) so tax paid is 40 on EBT of 180 = 40/180 = 22.22%

Leverage = Leverage ratio gives the the proportion of debt in comparison to equity = Debt / Equity (800/1200) = 66.67%

I/S 2015    2016
Sales (S) 1,200 1,320 (+10%)
- Costs (C) (COGS &  SG&A) 1,000 1100.00 COGS is increasing 10% as per the notes
= EBITDA (=EBIT) 200 220.00
- Interest 20 20
=EBT 180 200.00
- Tax (T) 40 44.44
=      NI 140 155.56
Dividend 40 44 Assuming Dividend increases by 10% every year
Plowback 100 111.56